A farmer grows a bushel of wheat and sells it to a miller for $1.00. The miller turns the wheat into flour and then sells the flour to a baker for $3.00. The baker uses the flour to make bread and sells the bread to an engineer for $6.00. The engineer eats the bread. What is the value added by each person? What is GDP?
1
Expert's answer
2020-05-11T19:41:43-0400

Each person adds:

The farmer: 1-0=1

The miller: 3-1=2

The baker: 6-3=3

GDP= final product cost=6 or GDP= amount of value added=1+2+3=6